OpenStudy (anonymous):

Multiplication of Rational Numbers #1: Solve... -7/9 x 3/5 #2: Solve... -2/5 x 1 1/4 #3: Solve... 1 3/5 x -3/4 #4: Solve... -7/8 x 1 3/7

OpenStudy (lstrasmann):

to multiply fractions you just have to multiply the numerators and denominators.

OpenStudy (anonymous):

@Istrasmann ...Not if the denominators don't match you can't. o_O

OpenStudy (lstrasmann):

no, the denominators only have to match for addition and subtraction

OpenStudy (anonymous):

@Istrasmann either way, I'm asking for the answers, not knowing how to do it. Otherwise then I'd look up how to do these.

OpenStudy (anonymous):

\[\frac{ -7 }{ 9}*\frac{ 3 }{ 5 } = \frac{ -7*3 }{ 9*5 } = \frac{ -21 }{ 45 }=\frac{ -7 }{15 }\]

OpenStudy (anonymous):

Like,,this.....@xclaireykinzx

OpenStudy (lstrasmann):

#1 is -21/45 #2 is 1 -1/10 #3 is 1 -9/20 #4 is 1 -21/56 And unless Mrs. Schaffnit was wrong, you do multiply even if the denominators are different. (She was my 7th grade teacher.)

OpenStudy (anonymous):

Do the Same for Other things too